OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (–8, 3); Slope: 6 y + 3 = 6x – 48 y + 3 = 6(x + 8) y – 3 = 6(x – 8) y – 3 = 6(x + 8)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

y -y1 = m(x-x1) where m =6

OpenStudy (anonymous):

(x1,y1) = (-8,3) put the values